# BloomScene: Lightweight Structured 3D Gaussian Splatting for Crossmodal Scene Generation

### v1

@@ Introduction
3D Gaussians Initialization. We use $\bm{P}_{N}$ as the initial SfM (Schonberger and Frahm 2016) points to initialize 3DGS. Since the initial $(N+1)$ views are not sufficient to train the 3DGS to produce reasonable outputs, we choose to add additional $M$ support views to form the image training set $\bm{I}_{i=0}^{N+M}$ of 3DGS. Unlike previous work (Chung et al. 2023), we choose to take the depth of the center of each depth map $\bm{D}_{i}$ as the radius of the spheres. The cameras are shifted ±$\pm 5^{\circ}$ along each sphere to get new cameras ${\{\bm{C}_{i}\}}_{i=N+1}^{N+M}$. The image training set $\bm{I}_{i=0}^{N+M}$ are obtained by reprojection from $\bm{P}_{N}$ using ${\{\bm{C}_{i}\}}_{i=0}^{N+M}$:
@@
$\mathcal{L}_{DPR}^{pixel}=\left\{\begin{array}[]{l}g_{\mathrm{rgb}}\frac{1}{|\hat{\bm{D}}|}\sum\|\bm{D}-\hat{\bm{D}}\|_{1},\text{ if }\|\bm{D}-\hat{\bm{D}}\|_{1}>\delta\\ g_{\mathrm{rgb}\mid}\frac{1}{|\bm{D}|}\sum\frac{(\bm{D}-\hat{\bm{D}})^{2}+\delta^{2}}{2\delta},\text{ otherwise }\end{array}\right.,$ (7)
where $g_{rgb}=exp(-\nabla)$ and $\nabla$ is the gradient of the current aligned RGB image, $\delta=0.2\max\|\bm{D}-\hat{\bm{D}}\|_{1}$, and $|\hat{\bm{D}}|$ indicates the total number of pixels in $\hat{\bm{D}}$. Image edges with larger gradients are dynamically assigned smaller learning weights. Constraining two depth maps only at the pixel level ignores the discrepancy of their distributions. Therefore, we implement distribution-level alignment between depth maps based on CMD, which has been widely used in domain adaptation to estimate the discrepancy between two domains (Zellinger et al. 2019). CMD can utilize higher-order moments to effectively capture higher-order statistical information without kernel function dependence. For a random variable $\bm{X}$, the $k$-th central moment is given by: $\mu_{k}=\mathbb{E}\left[(\bm{X}-\mathbb{E}[\bm{X}])^{k}\right]$, where $\mathbb{E}[\bm{X}]$ denotes the mean of $\bm{X}$. For two distributions $\bm{P}$ and $\bm{Q}$, the CMD computes the discrepancy by summing the differences of their corresponding central moments up to order $K$ :
@@
[…]
@@
where $\lambda_{4}$ and $\lambda_{5}$ are set to to $21e-$ 2 and $32e-$ 3, $\mathcal{L}_{vol}$ is the regularization term defined in (Lu et al. 2024).

### v2

@@ Introduction
3D Gaussians Initialization. We use $\bm{P}_{N}$ as the initial SfM (Schonberger and Frahm 2016) points to initialize 3DGS. Since the initial $(N+1)$ views are not sufficient to train the 3DGS to produce reasonable outputs, we choose to add additional $M$ support views to form the image training set $\bm{I}_{i=0}^{N+M}$ of 3DGS. Unlike previous work (Chung et al. 2023), we choose to take the depth of the center of each depth map $\bm{D}_{i}$ as the radius of the spheres. The cameras are shifted $\pm 5^{\circ}$ along each sphere to get new cameras ${\{\bm{C}_{i}\}}_{i=N+1}^{N+M}$. The image training set $\bm{I}_{i=0}^{N+M}$ are obtained by reprojection from $\bm{P}_{N}$ using ${\{\bm{C}_{i}\}}_{i=0}^{N+M}$:
@@
[…]
@@
$p\left(\hat{\bm{f}}_{i}\right)=\int_{\hat{\bm{f}}_{i}-\frac{1}{2}\omega_{i}}^{\hat{\bm{f}}_{i}+\frac{1}{2}\omega_{i}}\phi_{\mu_{i},\sigma_{i}}(x)dx,$ (14)
@@
where $\lambda_{4}$ and $\lambda_{5}$ are set to to 1e21e-2 and 2e32e-3, $\mathcal{L}_{vol}$ is the regularization term defined in (Lu et al. 2024).